Random variable X has the following probability distribution.

x 0 1 3 4 5 6

P .X D x/ 0.1 0.1 0.3 0.2 0.1 0.2

(a) Find P .3  X  5/. (3)

(b) Find E .X/: (5)

(c) Find V ar .X/ (7)


1
Expert's answer
2021-07-22T08:48:04-0400
"\def\arraystretch{1.5}
 \begin{array}{c:c}
 x & 0 & 1 & 3 & 4 & 5 & 6 \\ \hline
 P(X=x) & 0.1 & 0.1 & 0.3 & 0.2 & 0.1 & 0.2 \\

\end{array}"


Check


"0.1+0.1+0.3+0.2+0.1+0.2=1"

(a)


"P(3\leq X\leq 5)=P(X=3)+P(X=4)+P(X=5)"

"=0.3+0.2+0.1=0.6"

"P(3< X<5)=P(X=4)=0.2"



(b)


"E[X]=0(0.1)+1(0.1)+3(0.3)+4(0.2)"

"+5(0.1)+6(0.2)=3.5"



(c)


"E[X^2]=0^2(0.1)+1^2(0.1)+3^2(0.3)+4^2(0.2)"

"+5^2(0.1)+6^2(0.2)=15.7"




"Var(X)=E[X^2]-(E[X])^2"

"=15.7-(3.5)^2=3.45"